At a temperature of 32°C, the vapor pressure of water is 35.7 mmHg. If 28g of H2SO4 are added to 300g of H2O, what will the vapor pressure of the solution be?
1
Expert's answer
2020-04-27T01:46:44-0400

Relative lowering of vapour pressure is given by "\dfrac{p_o-p}{p_o}=mole\ fraction \ of \ solute"

"\dfrac{35.7-p}{35.7}=" "\dfrac{\frac{28}{98}}{\frac{28}{98}+\frac{300}{18}}=\dfrac{0.286}{0.286+16.67}=\dfrac{0.286}{16.956}"


"p=35.097mm \ Hg"